1.

Consider pdf of a Normal variable ๐‘ฟ is defined as ๐’‡(๐’™) = (1/โˆš98๐…)e^-((x+11)^2/98))

i) Find mgf of ๐‘ฟ.

ii) Evaluate ๐‘ท(โˆ’๐Ÿ‘ > โˆ’๐‘ฟ > ๐Ÿ๐Ÿ‘).

iii) Find value of ๐‘ช such that ๐‘ท(|๐‘ฟ + ๐Ÿ๐Ÿ| โ‰ฅ ๐‘ช) = ๐ŸŽ. ๐ŸŽ๐Ÿ”๐Ÿ๐Ÿ’.

iv) Find โˆ’๐’๐ŸŽ.๐ŸŽ๐ŸŽ๐Ÿ‘๐Ÿ and convert to ๐‘ฟ.


2.

Let ๐’€๐Ÿ < ๐’€๐Ÿ < ๐’€๐Ÿ‘ <๐’€๐Ÿ’ < ๐’€๐Ÿ“ be the order statistics of five independent observations ๐‘ฟ๐Ÿ, ๐‘ฟ๐Ÿ, ๐‘ฟ๐Ÿ‘, ๐‘ฟ๐Ÿ’ & ๐‘ฟ๐Ÿ“ each from distribution with pdf ๐’‡(๐’™) = ๐’†-2x; ๐ŸŽ โ‰ค ๐’™ < โˆž.

(i) Find pdf of the sample median and itโ€™s mean.

(ii) Determine ๐‘ท(๐’€๐Ÿ’ โ‰ฅ ๐ŸŽ. ๐Ÿ“).ย 

where, ๐‘ฎ๐’“ (๐’š) = \sumโˆ‘(nr ) [๐‘ญ(๐’š)] ๐’“ [๐Ÿ โˆ’ ๐‘ญ(๐’š)] ๐’โˆ’๐’“ ,

๐’ˆ๐’“ (๐’š) = (๐’!/(๐’โˆ’๐’“)!(๐’“โˆ’๐Ÿ)!) [๐‘ญ(๐’š)] ๐’“โˆ’๐Ÿ [๐Ÿ โˆ’ ๐‘ญ(๐’š)] ๐’โˆ’๐’“๐’‡(๐’š)ย 


3. Let ๐‘ฟ equals weight of a soap. A random sample of size ๐Ÿ๐Ÿ of ๐‘ฟ yielded with weights ๐Ÿ“๐Ÿ๐Ÿ‘, ๐Ÿ’๐Ÿ—๐Ÿ‘, ๐Ÿ’๐Ÿ—๐Ÿ–, ๐Ÿ“๐ŸŽ๐Ÿ‘, ๐Ÿ’๐Ÿ—๐Ÿ“, ๐Ÿ“๐Ÿ๐Ÿ, ๐Ÿ“๐ŸŽ๐Ÿ“, ๐Ÿ’๐Ÿ—๐Ÿ•, ๐Ÿ’๐Ÿ–๐Ÿ–, ๐Ÿ’๐Ÿ—๐Ÿ‘, ๐Ÿ“๐Ÿ๐Ÿ & ๐Ÿ“๐ŸŽ๐Ÿ‘

grams respectively. Now order them increasingly and find Median, ๐‘ซ๐Ÿ• and ๐‘ท๐Ÿ’๐Ÿ‘. Is

there any mode exists? Determine the semi-range of given weights.


1
Expert's answer
2021-10-03T17:46:03-0400

1.

i)

The moment generating function (MGF):

"m_X(t)=Ee^{tX}"

"m_X(t)=\int^{\infin}_{-\infin}e^{tx}f(x)dx=\int^{\infin}_{-\infin}\frac{e^{tx}e^{-(x+11)^2/98}}{\sqrt{98\pi}}dx="


"=\frac{e^{-121/98}}{\sqrt{98\pi}}\int^{\infin}_{-\infin}e^{-(x^2+x(98t+22))/98}dx=\frac{e^{-121/98}}{\sqrt{98\pi}}\cdot e^{(98t+22)^2/392}\sqrt{98\pi}=e^{(24.5t^2+11t)}"


ii)

"P(-3>-X>13)=\int^{-3}_{-\infin}f(-x)dx+\int^{\infin}_{13}f(-x)dx="


"=\int^{-3}_{-\infin}\frac{e^{-(-x+11)^2/98}}{\sqrt{98\pi}}dx+\int^{\infin}_{13}\frac{e^{-(-x+11)^2/98}}{\sqrt{98\pi}}dx=-\frac{(erf(14/\sqrt{98})-1)\sqrt{98\pi}}{2\sqrt{98\pi}}-\frac{(erf(2/\sqrt{98})-1)\sqrt{98\pi}}{2\sqrt{98\pi}}"

where erf is error function.

"erf(14/\sqrt{98})=0.95,\ erf(2/\sqrt{98})=0.22"

"P(-3>-X>13)=0.025+0.390=0.415"


iii)

"P(|X+11|\ge C)=2\int^{\infin}_{C-11}\frac{e^{-(x+11)^2/98}}{\sqrt{98\pi}}dx=0.0614"


"2\cdot-\frac{erf((\sqrt{2}(C-11)+11\sqrt{2})/14)-1}{2}=0.0614"

"erf((\sqrt{2}(C-11)+11\sqrt{2})/14)=1-0.0614=0.9386"

"(\sqrt{2}(C-11)+11\sqrt{2})/14=1.323"


"C=\frac{14\cdot 1.323-11\sqrt{2}}{\sqrt{2}}+11=13.1"


iv)

"Z_{0.0031}=-3.42"

"-Z_{0.0031}=3.42"

"\frac{x-\mu}{\sigma}=3.42"

"\mu=-11,\ \sigma=7"

"X=3.42\sigma+\mu=3.42\cdot7-11=12.94"


3.

488, 493, 493, 495, 497, 498, 503, 503, 505, 511, 511, 513

"median=(498+503)\/2=500.5"


Decile:

"D_i=i(N+1)\/10"

"D_7=7\cdot (12+1)/10=9.1"


Percentile:

"0.43N=0.43\cdot12=5.16"

"0.16\cdot(498-497)=0.16"

"P_{43}=5+0.16=5.16"


Modes are: 493, 503, 511


The semi-interquartile range isย one-half the difference between the first and third quartiles:

"IQR=(Q_3-Q_3)\/2=(511-493)\/2=9"


2.

i)

pdf of the sample median:

"q(x)=c_n[P(x_{med})(1-P(x_{med}))]^{(n-1)\/2}f(x)"

"P(x_{med})=0.5"

cnย is a coefficient that represents the number of ways a sample of (n-1)/2 values above median and (n-1)/2 below median can be arranged.

"c_n=2!=2"

"q(x)=2\cdot[0.5\cdot(1-0.5)]^{(5-1)/2}e^{-2x}=0.125e^{-2x}"


The expected value of the median of the sample is equal to the median of the distribution f(x).

"P(x_{med})=\int^{x_{med}}_0f(t)dt=\int^{x_{med}}_0e^{-2t}dt=0.5"


"-e^{-2t}\/2|^{x_{med}}_0=-e^{-2x_{med}}\/2+1\/2=0.5"

"E(X)=x_{med}=\infin"


ii)

"g_4(y)=\frac{5!}{3!}\cdot0.5^3\cdot0.5e^{-2y}=1.25e^{-2y}"


"P(Y_4\ge0.5)=\int^{\infin}_{0.5}g_4(y)dy=1.25\int^{\infin}_{0.5}e^{-2y}dy="


"=-1.25e^{-2y}/2|^{\infin}_{0.5}=1.25/(2e)=0.2299"